At a department store buying 3 belts and 4 ties cost 71 dollars,while 4 belts with 3 ties cost 69 dollars.Find the price of a belt and the price of a tie.

Answer by checkley79(3341)

3B+4T=71 MULTIPLY BY 4
4B+3T=69 MULTIPLY BY -3 & ADD.
12B+16T=284
-12B-9T=-207
---------------
7T=77
T=77/7
T=$11 FOR EACH TIE.
3B+4*11=71
3B+44=71
3B=71-44
3B=27
B=27/3
B=$9 OR EACH BELT.
PROOF:
4*9+3*11=69
36+33=69
69=69